OpenStudy (anonymous):

If a = 2 and b = -3, evaluate (a^2 - b^2)^3 A) -665 B) -125 C) -5 D) 2197

OpenStudy (anonymous):

B

zepdrix (zepdrix):

\[\huge \left[(2)^2-(-3)^2\right]^3\] \[\huge \left[4-9\right]^3\] \[\huge \left[-5\right]^3\]

OpenStudy (anonymous):

thanks to the both of yall.:}}